If a particle moves along a straight road with linearly increasing acceleration then what is
type of variation is exists between displacement and time?

Grade:8

1 Answers

Ashish Kumar
10 years ago
If acceleration is linearly incresing then , a= kt, where k is constant and t is the time.
a=dv/dt= kt, integrate both sides then v=kt^2/2.
v=dx/dt =kt^2/2. Integrate and x=kt^3/6.
so x is directly proportional to t^3.